OpenStudy (anonymous):

Parking in a student lot costs $1 for the first half hour and $1.50 for each thereafter. A partial hour is charged the same as a full hour. What is the longest time that a student can park in this lot for $10?

OpenStudy (anonymous):

i tried to solve it do you wana c mines

OpenStudy (anonymous):

$10-1=$9 9*1.50=13.5 round to 14 14*1.50= 21+1/2=23

OpenStudy (anonymous):

For this question, it's easier to solve if you set up an equation: so the total cost is $10 = $1+ $1.50(t) where t=# hours so solving this, you get $9 = $1.5t t = 6 hrs

OpenStudy (anonymous):

so then, the total time is 6.5 hrs

OpenStudy (anonymous):

because we've already accounted for the first half hour when we paid $1. Plus the 6 other hours

OpenStudy (anonymous):

what do you think

OpenStudy (anonymous):

what the heck i went wrong

OpenStudy (anonymous):

You multiplied 9 by 1.5 instead of dividing

OpenStudy (anonymous):

6.5 is 7 hours for the students to sit in parking lot

OpenStudy (anonymous):

no. $9/$1.5 per hr = 6 hrs. Add the original 1/2 hr that you gained from the $1 and you have 6.5 hrs. You don't need to round it

OpenStudy (anonymous):

oh ok well i see